# Systems Security Foundations for Agentic Computing

### v1

@@
Mihai Christodorescu1, Earlence Fernandes2, Ashish Hooda1,
Somesh Jha1,3, Johann Rehberger4, Khawaja Shams1
December 1, 2025
(May 2025)
This paper articulates short- and long-term research problems in AI agent security and privacy, using the lens of computer systems security. This approach examines end-to-end security properties of entire systems, rather than AI models in isolation. While we recognize that hardening a single model is useful, it is important to realize that it is often insufficient. By way of an analogy, creating a model that is always helpful and harmless is akin to creating software that is always helpful and harmless. The collective experience of decades of cybersecurity research and practice shows that this is insufficient. Rather, constructing an informed and realistic attacker model before building a system, applying hard-earned lessons from software security, and continuous improvement of security posture is a tried-and-tested approach to securing real computer systems. A key goal is to examine where research challenges arise when applying traditional security principles in the context of AI agents. A secondary goal of this report is to distill these ideas for AI and ML practitioners and researchers. We discuss the challenges of applying security principles to agentic computing, present 11 case studies of real attacks on agentic systems, and define a series of new research problems specific to the security of agentic systems.
@@ 1 Introduction
The IEEE Secure Generative AI Agents 2025 workshop (SAGAI; co-located with the IEEE Symposium on Security and Privacy) set out to explore how computer-security research and practice can help inform and advance agentic computing security. This report captures those discussions and distills key lessons, open research problems and recommendations on how the two fields can co-operate to build a foundation for secure agentic computing.
Agentic computing is when AI models interact with computer systems and services on behalf of human users by calling tools in a loop [willison2025agentdefinition]. The user specifies a task in natural language and the AI model is usually a large language/vision model but can be other types as well. The set of computer systems that can be operated by an agent include desktop interfaces, terminals, web browsers, mobile operating systems. Furthermore, agents can use arbitrary tools defined through various standards (e.g., REST or MCP [modelcontextprotocol_intro2025]).
Current efforts at securing agentic systems primary rely on model hardening approaches, often described as AI alignment [chen2025secalign, wallace2024instructionhierarchytrainingllms, openai2025safetyalignment]. Fundamentally, this involves training a model to resist attacks by exposing it to various exemplars. From a computer security perspective, this is equivalent to writing software that is always helpful, harmless and resistant to attacks. The past $50$ years of research and practice in computer security has empirically shown this to be an insufficient strategy. While a single program’s robustness should certainly be increased (e.g., computer security does this using formal methods approaches for certain kinds of programs or by writing software using languages (e.g. Rust) that are immune to certain classes of attacks), it has proven to be very difficult to scale to all types of computer programs. This has necessitated a systems approach that applies security defense techniques at various layers of abstraction of a computer system (i.e., hardware, firmware, hypervisors, operating systems, user-space frameworks, applications, network). The essence of this approach is to create deterministic guardrails at different layers of abstraction that increase the exploitation burden on the attacker while minimizing performance and functionality impact.
The situation with AI is worse. Models are not traditional computer programs that are immediately amenable to traditional computer-security techniques that increase their robustness. They are also vulnerable to inference-time attacks that allow attackers to control outputs in arbitrary ways [zou2023universaltransferableadversarialattacks, fu2024impromptertrickingllmagents, szegedy2014intriguingpropertiesneuralnetworks, greshake2023notwhatyouvesignedupfor]. Furthermore, the collective experience of the community between the two “waves” of adversarial machine learning research is that training models to resist attacks is insufficient: attackers always find workarounds [nasr2025attackermovessecondstronger, athalye2018obfuscatedgradientsfalsesense, pandya2025iattentionbreakingfinetuning]. Additionally, in the LLM-era of this research, attackers are not sophisticated and rely on English-language skills to trick AI systems into pursuing malicious tasks [greshake2023notwhatyouvesignedupfor, wunderwuzzi2024breaking, wunderwuzzi2025claude, wunderwuzzi2024chatgpt, wunderwuzzi2025devin]. Therefore, to make foundational and practical progress, the fields of computer security and AI agent safety/security need to co-operate.
Summary of Findings. Computer security relies on a range of security design principles to guide the exploration and implementation of deterministic guardrails [saltzer, ross-anderson]. Applying those principles to agentic systems encounters several technical challenges that we explain and explore in this report based on discussions during the workshop. Each of these challenges also pose interesting research challenges.
(1) Probabilistic TCB. A classic principle is creating a trusted computing base (TCB) to enforce a security invariant (e.g., to reduce the risk of memory corruption vulnerabilities, the hardware provides a deterministic feature that prevents code execution only from certain memory regions). Applying this to an agentic system is difficult because the model itself is a major part of the TCB, but it is fundamentally probabilistic. If you train the model to recognize that certain token sequences are non-executable, that property is enforced with some probability that is conditional on the input to the model rather than being a deterministic property. In other words, the challenge is to build provable defenses from probabilistic TCB assumptions.
(2) Dynamic Security Policies. A security policy governs the privilege of an untrusted program and is typically created by a developer or inferred based on code analysis. Applying this to an agentic system is difficult for two reasons: (a) there is no developer; only the end-user who specifies a natural language task and (b) there is no program to analyze; rather, the AI model “computes” many programs. Predicting or synthesizing a security policy given the natural language task is challenging because the task can be under-specified and using the agent’s entire context window for precise predictions is tricky because of the presence of untrusted external content.
(3) Fuzzy Security Boundary. Deterministic guardrails in computer security should be applied at the correct level of abstraction that provides the right amount of information to enforce the guardrail. For example, Android permissions are enforced at the interface between an application and the protected system services. Applying these principles to an agentic system is difficult because there aren’t any abstraction layers–typically the model just outputs a tool call with optional reasoning traces. Sometimes, these tool calls are too low-level to write and enforce reasonable security policies. For example, web agents perform UI-level manipulations making it difficult and brittle to write and enforce guardrails at the level of clicks and keystrokes. This is an instance of a well-known classic security problem of a semantic gap [bhushan-semantic-gap, brendan-semantic-gap, sommer-paxson].
@@
(4) Prompt Injections $\cong$ Dynamic Code Loading. Prompt injections are a key security problem to solve in agentic systems, but are sometimes necessary for functionality. For example, tool documentation does affect an agent’s behavior (by definition because the agent has to learn how to use the tool!). The challenge here is defining when prompt injections are harmful and when they are useful and then distinguishing the two. The analog in computer security is dynamic code loading–a very challenging problem that remains difficult to address today.
Scope and organization of this report. There has been a range of recent and contemporaneous writings on the general theme of applying computer security ideas to agentic computing [beurerkellner2025designpatternssecuringllm, zhang2025llmagentsemploysecurity]. The key distinction of this report is that we take a deep-dive on the technical challenges that arise when applying concrete security principles and articulate open research problems based on those technical challenges.
@@
The paper starts out with a discussion of classic computer-security design principles and security mechanisms (Section 2), intended for AI/ML practitioners and researchers. Following that, we discuss the application of these security principles to agentic settings and surface the challenges that arise forthwith (Section 3). We introduce in Section 4 case studies of real attacks on agentic systems with three goals in mind: (1) to showcase how real attacks are carried out by adversaries without the need for sophisticated technical tools to execute them; (2) to highlight the key security principle(s) that were violated; and (3) to outline the concrete security mechanisms that stop each such attack. These case studies not only illustrate how existing security mechanisms go a long way towards solving present-day attacks on agentic systems, but also show the types of the gaps that appear due to the challenges discussed earlier. Closing these gaps requires new research to advance the state of the art in agentic security—we summarize the open research problems in Section 5.
Figure 1: Standard security architecture showing trusted components in blue and untrusted ones in red.
2 Security Principles and Mechanisms
The standard security architecture consists of the trusted computing base (TCB), a security policy, the security boundary, and the untrusted system, as shown in Figure 1. The TCB consists of the functionality (code and data) whose integrity and confidentiality cannot be impacted by an attacker—in other words, the TCB defines the parts of the overall system that can be trusted to operate correctly under attack. The TCB typically contains the core functionality of the system (e.g., the kernel of the operating system) as well as a reference monitor that serves to approve or reject each request from the untrusted components to the TCB. The reference monitor is the part of the TCB that examines the security policy in order to make the approve/reject determination. The security policy is typically a declarative expression of the security goals—as an example, the security policy for stored files is given as a list of access control entries, each entry representing the operation(s) a given user is allowed to perform on a given file. The security boundary represents the interface through which the untrusted components interact with the TCB. In an operating system (OS), the security boundary between the OS kernel and the (untrusted) applications is the system-call interface.
Within this context we can define and describe the security principles introduced in the case studies of Section 4.
Principle of Least Privilege. This principle dictates that the Security Policy should only grant the Untrusted System the absolute minimum permissions it needs to function. For example, if the Untrusted System only needs to read a specific piece of data, the Security Policy should explicitly deny it permission to write, delete, or access any other data. The TCB is responsible for enforcing this minimal set of permissions, as specified by the Security Policy. For convenience most policies are written with the expectation of a default-deny fallback, where access to a resource not explicitly listed in the policy is automatically rejected.
TCB Tamper Resistance. This principle states that the TCB itself must be protected from modification by any outside influence. The Security Boundary must be designed to prevent the Untrusted System from altering the code or logic of the TCB or the Security Policy. If the TCB could be tampered with, an attacker could disable the “Check for permission” and bypass all security controls.
Complete Mediation. This principle dictates that every single request from the Untrusted System must be validated against the Security Policy. The system must not “remember” a previous authorization as users, trust levels, or other contextual conditions may have changed. Each time one of the requests crosses the Security Boundary, it must be intercepted by the TCB, which in turn must perform the “Check for permission to answer request” operation against the Security Policy. No request is allowed to bypass this check.
Secure Information Flow. This principle governs the “Requests & Responses” channel, ensuring that sensitive information does not leak to untrusted areas. The Security Policy must define what kind of information is allowed to flow in which direction. For instance, even if a request is valid, the TCB must check the Security Policy to ensure that the resulting Response does not contain secret data that the Untrusted System is not authorized to see (this would constitute an unauthorized flow of data from high-trust components to low-trust components).
Human Weak Link. This principle states that human operators can compromise this architecture in several ways, and thus the security mechanisms must be designed with this in mind. As a user, a human operating the Untrusted System could send malicious requests (e.g., SQL injection, prompt injection). As an administrator, a human could incorrectly configure the Security Policy, making it too permissive (violating Least Privilege). As a developer, a human could accidentally introduce a bug into the TCB that fails to perform the Check for permission operation correctly (violating Complete Mediation).
@@ 3.1 The TCB Is Probabilistic
In contrast, the TCB in agentic systems is built on LLMs that do not guarantee deterministic behavior—the same prompt can yield different outputs due to sampling randomness. Thus a fundamental challenge in realizing this traditional security principle in agentic systems is the TCB being probabilistic or non-deterministic (e.g. imagine whether we could build a memory safety defense on top of a probabilistic NX bit). This probabilistic nature undermines traditional security principles in three ways. First, a probabilistic reference monitor might correctly deny unauthorized access 99% of the time, but the remaining 1% can be exploited. Second, unlike traditional systems where security properties can be formally verified, we cannot prove an LLM will always enforce policies correctly. Third, LLMs operate in continuous representation spaces, making them vulnerable to adversarial examples. Unlike traditional systems that have discrete instruction sets with clear boundaries, attackers can use various search and optimization techniques to efficiently find inputs that appear benign but cause policy violations.
To build secure systems on probabilistic LLMs, we must leverages external deterministic information to disambiguate LLM decisions. An example of a system takes this approach is SkillFence [10.1145/3517232]. It uses deterministic signals from the user’s browsing history and installed apps to correct any mistakes made by a voice assistant’s language understanding. However, we believe that building provable defenses on probabilistic and non-deterministic TCB is a challenging—perhaps impossible—task.
@@ 3.2 The Security Policy Is Dynamic and Task-Specific
    The agent can navigate to GitLab issue 7745;
@@
Now, how should the privilege of the agent change to solve the new task description? One option is to reset the agent’s context and redo the policy prediction. This approach is secure but can result in lost utility because the agent’s context is erased. Thus, how can we evolve the security policy in a secure manner while there is untrusted data in the context? We believe that the key lies in creating policy languages that are amenable to formalized reasoning so that when new policies are predicted and added to the existing set, we can reason about how the new policies change the overall privilege level of the agent.
@@ 3.3 The Security Boundary Is Fuzzy
Identifying the right abstraction at which security policies are enforced is crucial to upholding the complete mediation design principle. In traditional computer systems, we have layered architectures (Hardware - OS Kernel - Process - Network) that offer different levels of abstraction for enforcing security policies. For example, the process-kernel interface allows for SELinux-style MAC policies and the App-Runtime interface allows for higher-level permissions that we see in systems like Android. Agentic systems lack these layers. Rather, they directly use tools given a natural language prompt. There are no layers in between. Thus, enforcing a security policy at the final layer of tool-calls can lead to incomplete or ineffective mediation (i.e., there can be gaps). The key research challenge in applying the complete mediation principle is identifying or creating the right abstraction at which a security policy can be enforced.
@@
CaMeL [camel] and FIDES [fides] introduce an agent design that exemplifies the “solution as code” approach. To solve a task, the CaMeL/FIDES agent first generates code that, when executed, would solve the task specified in natural language. Notice that this introduces a new layer of abstraction: the code needed to solve the task. At that point, one can analyze the code using standard security analysis methods (control and data flows) to determine whether the agent’s plan is aligned with the user’s goal. A key assumption in this design is that the agent has access to a set of semantically-meaningful tools. One of the key future research questions here is creating an appropriate intermediate representation for agent plans that is amenable to static and dynamic analysis.
@@
ceLLMate [cellmate] introduces a sandboxing framework for browser agents. The key challenge is identifying the right abstraction to enforce security policies within existing infrastructure. Browser agents, as currently built, have access to low-level UI manipulation tools (e.g., clicks and keystrokes), and thus, it is non-sensical to write policies at that level. The key insight to address this semantic gap is realizing that no matter what UI manipulations the agents perform, HTTP-level messages capture those actions in semantically meaningful ways. Therefore, ceLLMate policies are specified and enforced at the HTTP level, in co-operation with information supplied by webserver developers. One of the key future research challenges here is automatically predicting policies at the level of HTTP messages using natural language task descriptions as input.
3.4 Prompt Injection $\cong$ Dynamic Code Loading
Prompt injection, more specifically indirect prompt injection, is seen as a vulnerability of current AI-based systems. But prompt injection is a manifestation of a new class of functionality in AI, where the agent can adjust its task based on new information in its context. For example, using the GitLab task in Section 3.2, the agent may start with this prompt:

### v2

# spacing=nonfrench spacing=nonfrench

@@
Mihai Christodorescu1  Earlence Fernandes2  Ashish Hooda1  Somesh Jha1,3
Johann Rehberger4  Kamalika Chaudhuri5  Xiaohan Fu2,6  Khawaja Shams1
Guy Amir7  Jihye Choi3  Sarthak Choudhary3  Nils Palumbo3
Andrey Labunets2  Nishit V. Pandya2
1Google    2University of California San Diego    3University of Wisconsin–Madison
4EmbraceTheRed    5FAIR at Meta    6Gray Swan AI    7Cornell University
In recent years, agentic artificial intelligence (AI) systems are becoming increasingly widespread.
These systems allow agents to use various tools, such as web browsers, compilers, and more. However, despite their popularity, agentic AI systems also introduce a myriad of security concerns, due to their constant interaction with third-party servers. For example, a malicious adversary can cause data exfiltration by executing prompt injection attacks, as well as other unwarranted behavior. These security concerns have recently motivated researchers to improve the safety and reliability of agentic systems. However, most of the literature on this topic is from the AI standpoint and lacks the system-security perspective and guarantees. In this work, we begin bridging this gap and present an analysis through the lens of classic cybersecurity research. Specifically, motivated by decades of progress in this domain, we identify short- and long-term research problems in agentic AI safety by examining end-to-end security properties of entire systems, rather than standalone AI models running in isolation. Our key goal is to examine where research challenges arise when applying traditional security principles in the context of AI agents and, as a secondary goal, distill these ideas for AI practitioners. Furthermore, we extensively cover 11 case studies of real-world attacks on agentic systems, as well as define a series of new research problems that are specific to this important domain.
@@ 1 Introduction
Agentic AI systems have become widespread: by being given access to tools such as web browsers, compilers, and more, AI agents, and especially those based on large language models (LLM), are becoming increasingly powerful. However, this comes at a cost: access to third party tools and external servers introduces a wide range of security challenges that must be addressed in order to safely and reliably deploy such systems in a trustworthy manner. Previous literature has raised the important topic of agentic AI security, however, most of the this literature is written from an AI perspective, and does not leverage the extensive research accumulating through dozens of years of classic security studies. In this Systematization of Knowledge (SoK) paper, we begin bridging this gap, and explore how to build upon well-studied concepts and practices in computer-security research to advance the security of agentic computing. We focus on key similarities and differences between both domains, and distill lessons, open research problems and recommendations on how the two fields can co-operate to build a foundation for secure agentic computing.
Agentic computing refers to AI models that interact with computer systems and services on behalf of human users by repeatedly calling tools in a loop [100]. The user specifies a task in natural language, and the AI model (typically a large language or vision model) completes it by iteratively invoking tools, such as desktop UIs, web browsers, or mobile operating systems. For instance, a user might ask an agent to find mentions of a particular event in recent emails, summarize what happened, and save the summary to a text file. Such a request can trigger tools calls including email search, retrieval, and writes to local storage. Recently, there is also a trend in promoting uniform standards and protocols for defining various agentic tool calls, e.g., REST or Model Context Protocol (MCP) [58].
Current efforts to secure agentic systems span multiple layers of defense [8], including model-level, system-level, and user-level approaches. At the model level, various methods aim to make the agent more resilient to attacks by hardening approaches, often described as alignment techniques [18, 96, 62]. At the system level, an increasingly prominent line of work treats the model as an untrusted (or at best, probabilistic) component and instead enforces invariants through isolation and mediation: restricting the agent’s action space, sandboxing tool execution, structuring tool interfaces, and separating untrusted data from control flow via explicit trust boundaries and guardrails [8, 110]. Finally, at the user level, practical deployments commonly incorporate governance mechanisms such as explicit approvals/confirmations for sensitive actions, permission scoping, and audit/monitoring to reduce the blast radius of failures and support incident response [8].
From a computer-security perspective, this emerging taxonomy mirrors a classic defense-in-depth lesson: hardening a component is valuable, but is rarely sufficient on its own. Robust systems rely on deterministic enforcement points (e.g., access control, isolation, and provenance boundaries) placed at multiple layers of abstraction (hardware, hypervisors, operating systems, user-space frameworks, applications, and networks), to increase the attacker’s burden while preserving performance and functionality [82, 4]. In agentic settings, however, deciding where to place such enforcement points is less straightforward, because the model simultaneously (i) processes untrusted content and (ii) synthesizes actions and tool calls on the fly.
This tension is aggravated further due to modern agents being vulnerable to inference-time attacks that can steer behavior in arbitrary ways [115, 31, 91, 39]. As with earlier waves of adversarial ML, purely behavioral defenses that rely on the model to reliably identify and ignore malicious instructions can remain brittle under adaptive attackers [60, 7, 65]. Moreover, in the LLM era, attackers often need not have deep system expertise: simple prompting can be enough to induce harmful tool use or policy violations in real-world deployments [39, 70, 79, 74, 80]. Therefore, to make foundational and practical progress, the fields of computer security and AI agent security need to co-operate: combining model-level robustness with principled system-level and user-level guardrails grounded in decades of security research.
Summary of Findings. Computer security makes progress by articulating clear security invariants and then enforcing them via layered, largely deterministic mechanisms and design principles (e.g., least privilege, complete mediation, separation of privilege) [82, 4]. When we attempt to apply these ideas to agentic systems, we encounter several challenges that arise. We use these frictions to organize the report and to surface concrete research questions. Furthermore, each of these challenges also poses interesting research avenues for long-term future work.
(1) Probabilistic and Opaque TCB. A classic systems-security move is to minimize a trusted computing base (TCB) that deterministically enforces an invariant (e.g., preventing execution from non-executable memory regions to reduce the impact of memory-corruption bugs). In many AI-based systems, including agentic ones, the model itself becomes a central part of the TCB, yet its behavior is fundamentally probabilistic and often not fully interpretable. If we “train the model to never do XX”, that guarantee is inherently conditional and statistical: it holds with some probability that depends on the input distribution and the surrounding context rather than as an absolute property of the system. The resulting challenge is to build defenses with meaningful assurance guarantees despite relying on a probabilistic/opaque component in the TCB.
(2) Dynamic Security Policies. A security policy specifies what an untrusted program is allowed to do and is typically authored by a developer or inferred from code analysis. However, in agentic settings, both these assumptions break down: (i) there may be no developer intent beyond the end-user’s natural-language request, and (ii) there is no fixed program to analyze, since the agent effectively synthesizes many programs on the fly. Inferring or synthesizing a policy from a task description is difficult because tasks are often underspecified, and using the agent’s full context window for precise policy inference is nontrivial due to untrusted external content.
(3) Fuzzy Security Boundary. Deterministic guardrails are most effective when enforced at an interface that exposes the right semantic information (e.g., Android permissions at the boundary between apps and protected services). In many agentic stacks, however, the boundary between “decision” and “action” is blurry: the model may directly emit tool calls (sometimes via low-level UI actions) without stable abstraction layers that cleanly separate intent from mechanism. As a result, policies written at too low a level (clicks/keystrokes) become brittle, while policies written at too high a level may lack enforceable grounding. This can also be viewed as a manifestation of the classic semantic gap problem [44, 29, 89].
@@
(4) Dynamic Instruction Following. Prompt injections are a key security problem to solve in agentic systems, but the underlying mechanism of dynamically updating agentic instructions is oftentimes necessary for functionality and thus not necessarily malicious. For example, tool documentation does affect an agent’s behavior (by definition because the agent has to learn how to use the tool!). The challenge here is defining when prompt injections are harmful [26] and when they are useful and then distinguishing the two. The analog in computer security is dynamic code loading–a very challenging problem that remains difficult to address today.
@@
Scope and organization of this report. There has been a range of recent and contemporaneous writings on the general theme of applying computer security ideas to agentic computing [8, 110]. The key distinction of this report is that we take a deep-dive on the technical challenges that arise when applying concrete security principles and articulate open research problems based on those technical challenges. The paper starts out with a discussion of security principles applied to agentic settings and the challenges that arise forthwith (Section 2). Section 3 presents 11 case studies of real attacks on agentic systems with three goals in mind: (1) to showcase how real attacks are carried out by adversaries without the need for sophisticated technical tools to execute them; (2) to highlight the corresponding security principle(s) that were violated; and (3) to outline concrete security mechanisms needed to stop each such attack. We discuss the current practices for mitigating these threats in Section 4 and summarize the open research problems in Section 5.
@@ 2.1 The TCB is Probabilistic
In contrast, the TCB in agentic systems is built on LLMs that do not guarantee predictable, or even deterministic behavior—the same prompt can yield different outputs due to sampling randomness. Thus a fundamental challenge in realizing this traditional security principle in agentic systems is the TCB being probabilistic or non-deterministic (e.g. imagine whether we could build a memory safety defense on top of a probabilistic NX bit). This probabilistic nature undermines traditional security principles in three ways. First, a probabilistic reference monitor might correctly deny unauthorized access 99% of the time, but the remaining 1% can be exploited. Second, unlike traditional systems where security properties can be formally verified, we cannot prove an LLM will always enforce policies correctly. Third, LLMs operate in continuous representation spaces, making them vulnerable to adversarial examples. Unlike traditional systems that have discrete instruction sets with clear boundaries, attackers can use various search and optimization techniques to efficiently find inputs that appear benign but cause policy violations.
To build secure systems on probabilistic LLMs, we must leverage external deterministic information to disambiguate LLM decisions. An example of a system takes this approach is SkillFence [41]. It uses deterministic signals from the user’s browsing history and installed apps to correct any mistakes made by a voice assistant’s language understanding. However, we believe that building provable defenses on probabilistic and non-deterministic TCB is a challenging—perhaps impossible—task.
@@ 2.2 The Security Policy Is Dynamic and Task-Specific
    The agent can navigate to GitLab issue 745;
@@
Now, how should the privilege of the agent change to solve the new task description? One option is to reset the agent’s context and redo the policy prediction. This approach is secure but can result in lost utility because the agent’s context is erased. Thus, how can we evolve the security policy in a secure manner while there is untrusted data in the context?
An additional problem for AI agents is that privacy consent is fundamentally at odds with agency  [93], as finding and enforcing the right policy for an AI agent each time requires resolving the trade-off between providing more access for better task performance and restricting access as per least-privilege principle. We believe that the key lies in creating policy languages that are amenable to formalized reasoning so that when new policies are predicted and added to the existing set, we can reason about how the new policies change the overall privilege level of the agent.
@@ 2.3 The Security Boundary Is Fuzzy
Identifying the right abstraction at which security policies are enforced is crucial to upholding the complete mediation design principle. In traditional computer systems, we have layered architectures (Hardware—Operating System (OS) Kernel—Process—Network) that offer different levels of abstraction for enforcing security policies. For example, the process-kernel interface allows for SELinux-style MAC policies and the App-Runtime interface allows for higher-level permissions that we see in systems like Android. Agentic systems lack these layers. Rather, they directly use tools given a natural language prompt. There are no layers in between. Thus, enforcing a security policy at the final layer of tool-calls can lead to incomplete or ineffective mediation (i.e., there can be gaps). The key research challenge in applying the complete mediation principle is identifying or creating the right abstraction at which a security policy can be enforced.
@@
CaMeL [25] and FIDES [23] introduce an agent design that exemplifies the “solution as code” approach. To solve a task, the agent first generates code that, when executed, would solve the task specified in natural language. Notice that this introduces a new layer of abstraction: the code needed to solve the task. At that point, one can analyze the code using standard security analysis methods (control and data flows) to determine whether the agent’s plan is aligned with the user’s goal. A key assumption in this design is that the agent has access to a set of semantically-meaningful tools. One of the key future research questions here is creating an appropriate intermediate representation for agent plans that is amenable to static and dynamic analysis.
@@
ceLLMate [54] introduces a sandboxing framework for browser agents. The key challenge is identifying the right abstraction to enforce security policies within existing infrastructure. Browser agents, as currently built, have access to low-level UI manipulation tools (e.g., clicks and keystrokes), and thus, it is non-sensical to write policies at that level. The key insight to address this semantic gap is realizing that no matter what UI manipulations the agents perform, HTTP-level messages capture those actions in semantically meaningful ways. Therefore, ceLLMate policies are specified and enforced at the HTTP level, in co-operation with information supplied by webserver developers. One of the key future research challenges here is automatically predicting policies at the level of HTTP messages using natural language task descriptions as input.
2.4 Dynamic Instruction Following